# The isoperimetric inequality for the capillary energy outside convex sets

### v1

@@ Abstract.
We study the isoperimetric problem for capillary hypersurfaces with a general contact angle $\theta\in(0,\pi)$, outside arbitrary convex sets. We prove that the capillary energy of any surface supported on any such convex set is larger than that of a spherical cap with the same volume and the same contact angle on a flat support, and we characterize the equality cases. This provides a complete solution to the isoperimetric problem for capillary surfaces outside convex sets at arbitrary contact angles, generalizing the well-known Choe–Ghomi–Ritoré inequality, which corresponds to the case $\theta=\frac{\pi}{2}$.
@@ 1.1. Overview of the proof
The proof of Theorem 1.1 is based on the ABP-method applied to the Neumann problem (1.3). We note that in the context of isoperimetric problems this method was used first time by Cabré in  [3, 4] and further generalized in [5, 9]. A variant of this method has been used recently in [23] to prove the stability of the isoperimetric inequality for the capillary energy in a half space (see also the recent preprint [6] for a stronger quantitative version of the same isoperimetric inequality).
@@
$\begin{cases}&u=c\quad\text{in }\,\Omega\\ &\partial_{\nu}u=1\quad\text{on }\,\Sigma\\ &\partial_{\nu}u=-\lambda\quad\text{on }\,\Gamma,\end{cases}$ (1.3)
@@
In order to deal with the case of general $\lambda$’s, we proceed as follows. We first show that the (variational) solution to (1.3) is a viscosity supersolution to the same problem, in the sense of Definition 2.1. Note that the latter definition is stronger than the one given in  [11] and therefore the supersolution property in the above sense does not follow from known results. Using this property we are able to relate the subdifferentials of $u$ in $\Omega$ with the subdifferentials of the restriction of $u$ to $\Gamma$ proving the following inclusion, see Lemma 2.4,
@@ 2.2. Subdifferentials and restricted subdifferentials
By the above it holds that the inequality $u(y)\geq\varphi(y)$ is not true for all $y\in\overline{\Omega}$. This means that $\bar{c}=\min_{y\in\overline{\Omega}}\,(u(y)-\varphi(y))<0$. In turn we have that the graph of $\varphi+\bar{c}$ touches from below the graph of $u$ at some point $\bar{x}\in\overline{\Omega}$. Clearly $\bar{x}\not\in\Gamma$, since $\xi\in J_{\Gamma}u(x)$ and $\bar{c}<0$ imply $\varphi(y)+\bar{c}<u(y)$ for all $y\in\Gamma$. On the other hand, if $\bar{x}\in\Sigma$, again by the supersolution property, we would have that $\xi\cdot\nu_{\Sigma}(\bar{x})\geq 1$, which is impossible as $|\xi|<1$. Therefore $\bar{x}\in\Omega$, which implies $\xi=\nabla u(\bar{x})$, as desired. ∎
@@ 3. The main geometric estimate
The next useful observation is that by discretization we can reduce the proof of (3.1) to the case of functions $v$ defined on finite subsets of $\partial\mathbf{C}$, as shown in the following lemma. To this aim we recall that a sequence $\{C_{n}\}$ of closed sets of $\mathbb{R}^{N}$ converge in the Kuratoswki sense to a closed set $C$ if the following conditions are satisfied:
@@
Combining the previous proof with the ABP argument skeched in the introduction and rigorously developed in the next section, we recover an ABP-proof of the relative isoperimetric inequality outside convex sets obtained by Choe, Ghomi and Ritoré in [8]. Note that an ABP-argument for the same inequality has been already provided in [18]. However, in their argument our crucial estimate (3.9) is replaced by a geometric estimate based on normal cones and inspired by the techniques in [7], see [18, Proposition 2.4].
@@
Arguing as before, by the area and layer cake formulas, we get
@@ 4. The capillary isoperimetric inequality outside convex sets
and, in turn, since $\Delta u=N$,
@@ 5. Appendix
To approximate $E$ we first extend $\hbox{{\large$\chi$}{\Large${}_{{}_{E}}$}}_{\big{|}_{\mathbb{R}^{N}\setminus\mathbf{C}}}$ to a function $u\in BV(\mathbb{R}^{N})$, with compact support, such that $|Du|(\partial\mathbf{C})=0$, $0\leq u\leq 1$, see [2, Proposition 3.21]. Note that for all $t\in(0,1)$, $\{u>t\}\setminus\mathbf{C}=E$. For any $\varepsilon>0,t\in(0,1)$ we set $U_{\varepsilon,t}=\{x:\,u_{\varepsilon}(x)>t\}$, where $u_{\varepsilon}=\varrho_{\varepsilon}*u$, for a standard mollifier $\varrho_{\varepsilon}$. Note that for a.e. $t\in(0,1)$ there exists a sequence $\varepsilon_{n}$ converging to zero such that and
@@
Finally, property (vi) follows by observing that if $x\in\partial\Omega\setminus\Sigma_{sing}$ and ${\rm sd}_{\Omega}$ is smooth in $\overline{B}_{2r}(x)$, then ${\rm sd}^{\varepsilon_{m_{j}}}_{\Omega}\to{\rm sd}_{\Omega}$ in $C^{\infty}(\overline{B}_{r}(x))$, which in turn easily implies (vi). ∎
@@
N. F., M. M. and A. P. have been supported by PRIN 2022 Project “Geometric Evolution Problems and Shape Optimization (GEPSO)”, PNRR Italia Domani, financed by European Union via the Program NextGenerationEU, CUP D53D23005820006. N. F. and M. M. are members of the Gruppo Nazionale per l’Analisi Matematica, la Probabilità e le loro Applicazioni (GNAMPA), which is part of the Istituto Nazionale di Alta Matematica (INdAM). V. J. was supported by the Academy of Finland grant 314227.

### v2

@@ Abstract.
We study the isoperimetric problem for capillary hypersurfaces with a general contact angle $\theta\in(0,\pi)$, outside arbitrary convex sets. We prove that the capillary energy of any surface supported on any such convex set is larger than that of a spherical cap with the same volume and the same contact angle on a flat support, and we characterize the equality cases. This provides a complete solution to the isoperimetric problem for capillary surfaces outside convex sets at arbitrary contact angles, generalizing the well-known Choe-Ghomi-Ritoré inequality, which corresponds to the case $\theta=\frac{\pi}{2}$.
@@ 1.1. Overview of the proof
The proof of Theorem 1.1 is based on the ABP-method applied to the Neumann problem (1.3). We note that in the context of isoperimetric problems this method was used first time by Cabré in  [3, 4] and further generalized in [5, 9]. A variant of this method has been used recently in [24] to prove the stability of the isoperimetric inequality for the capillary energy in a half space (see also [17] and the recent preprint [6] for a stronger quantitative version of the same isoperimetric inequality).
@@
$\begin{cases}&\Delta u=c\quad\text{in }\,\Omega\\ &\partial_{\nu}u=1\quad\text{on }\,\Sigma\\ &\partial_{\nu}u=-\lambda\quad\text{on }\,\Gamma,\end{cases}$ (1.3)
@@
In order to deal with the case of general $\lambda$’s, we proceed as follows. We first show that the (variational) solution to (1.3) is a viscosity supersolution to the same problem, in the sense of Definition 2.1. The latter definition is stronger than the one given in  [11] and therefore the supersolution property in the above sense does not follow from known results. Using this property we are able to relate the subdifferentials of $u$ in $\Omega$ with the subdifferentials of the restriction of $u$ to $\Gamma$ proving the following inclusion, see Lemma 2.4,
@@ 2.2. Subdifferentials and restricted subdifferentials
By the above it holds that the inequality $u(y)\geq\varphi(y)$ is not true for all $y\in\overline{\Omega}$. This means that $\bar{c}=\min_{y\in\overline{\Omega}}\,(u(y)-\varphi(y))<0$. In turn we have that the graph of $\varphi+\bar{c}$ touches from below the graph of $u$ at some point $\bar{x}\in\overline{\Omega}$. Clearly $\bar{x}\not\in\Gamma$, since $\xi\in J_{\Gamma}u(x_{0})$ and $\bar{c}<0$ imply $\varphi(y)+\bar{c}<u(y)$ for all $y\in\Gamma$. On the other hand, if $\bar{x}\in\Sigma$, again by the supersolution property, we would have that $\xi\cdot\nu_{\Sigma}(\bar{x})\geq 1$, which is impossible as $|\xi|<1$. Therefore $\bar{x}\in\Omega$, which implies $\xi=\nabla u(\bar{x})$, as desired. ∎
@@ 3. The main geometric estimate
The next useful observation is that by discretization we can reduce the proof of (3.1) to the case of functions $v$ defined on finite subsets of $\partial\mathbf{C}$, as shown in the following lemma. To this aim we recall that a sequence $\{C_{n}\}$ of closed sets of $\mathbb{R}^{N}$ converges in the Kuratoswki sense to a closed set $C$ if the following conditions are satisfied:
@@
Combining the previous proof with the ABP argument skeched in the introduction and rigorously developed in the next section, we recover an ABP-proof of the relative isoperimetric inequality outside convex sets obtained by Choe, Ghomi and Ritoré in [8]. Note that an ABP-argument for the same inequality has been already provided in [19]. However, in their argument our crucial estimate (3.9) is replaced by a geometric estimate based on normal cones and inspired by the techniques in [7], see [19, Proposition 2.4].
@@
Arguing as before, by the Area and layer-cake formulas, we get
@@ 4. The capillary isoperimetric inequality outside convex sets
and, in turn, since $\det\nabla^{2}u\leq\big(\frac{\Delta u}{N}\big)^{N}=1$,
@@ 5. Appendix
To approximate $E$ we first extend $\hbox{{\large$\chi$}{\Large${}_{{}_{E}}$}}_{\big|_{\mathbb{R}^{N}\setminus\mathbf{C}}}$ to a function $u\in BV(\mathbb{R}^{N})$, with compact support, such that $|Du|(\partial\mathbf{C})=0$, $0\leq u\leq 1$, see [2, Proposition 3.21]. Note that for all $t\in(0,1)$, $\{u>t\}\setminus\mathbf{C}=E$. For any $\varepsilon>0,t\in(0,1)$ we set $U_{\varepsilon,t}=\{x:\,u_{\varepsilon}(x)>t\}$, where $u_{\varepsilon}=\varrho_{\varepsilon}*u$, for a standard mollifier $\varrho_{\varepsilon}$. Note that for a.e. $t\in(0,1)$ there exists a sequence $\varepsilon_{n}$ converging to zero such that
@@
Finally, property (vi) follows by observing that if $x\in\partial\Omega\setminus\Sigma_{sing}$ and ${\rm sd}_{\Omega}$ is smooth in $\overline{B}_{2\rho}(x)$, then ${\rm sd}^{\varepsilon_{m_{j}}}_{\Omega}\to{\rm sd}_{\Omega}$ in $C^{\infty}(\overline{B}_{\rho}(x))$, which in turn easily implies (vi). ∎
@@
N. F., M. M. and A. P. were supported by PRIN Project 2022E9CF89, “Geometric Evolution Problems and Shape Optimization (GEPSO)”, PNRR Italia Domani, financed by European Union via the Program NextGenerationEU. N. F., M. M. and A. P. are members of the Gruppo Nazionale per l’Analisi Matematica, la Probabilità e le loro Applicazioni (GNAMPA), which is part of the Istituto Nazionale di Alta Matematica (INdAM). V. J. was supported by the Academy of Finland grant 314227.